OpenStudy (anonymous):

f(x) = cos(sin(1/x)). Is the range of this function still [-1,1]? If so, why? f(x) = cos(sin(1/x)). Is the range of this function still [-1,1]? If so, why? @Mathematics

OpenStudy (jamesj):

First of all, what's the range of the function g(x) = sin(1/x)?

OpenStudy (jamesj):

We need to know that so we know the domain of the function f(u) = cos(u), where u = g(x).